# Adversarial Examples are Misaligned in Diffusion Model Manifolds

### 17 Jan 2024 v3

@@ 1 INTRODUCTION
Figure 1: Illustration of the difference between an adversarial and a benign sample when subjected to the transformation process. $p_{a}({\bm{x}})$ represents the distribution of adversarial images, while $p_{b}({\bm{x}})$ represents the distribution of benign images. Using the inversion and reversion process of DDIM [19], ${\bm{x}}_{a}$ and ${\bm{x}}_{b}$ become ${\bm{x}}_{a}^{\prime}$ and ${\bm{x}}_{b}^{\prime}$, respectively. These transformed counterparts now belong to distinct distributions, namely $p_{a}^{\prime}({\bm{x}})$ and $p_{b}^{\prime}({\bm{x}})$, characterized by a significantly reduced overlap. Therefore, this results in a distinct representation of adversarial samples compared to benign samples.
@@ 3 METHOD
In this section, we begin with reviewing DDPMs, and the inverse and reverse process of the DDIM [19]. Then, we present details of our method and how to capture the adversarial attack fingerprints.
@@ 3.1 Prelimaries
$\small q({\bm{x}}_{t}|{\bm{x}}_{t-1})=\mathcal{N}\left({\bm{x}}_{t};\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}},{\bm{x}}_{t-1},(1-\frac{\alpha_{t}}{\alpha_{t-1}})\mathbf{I}\right),$ (5)
in which ${\bm{x}}_{t}$ represents the noisy image at the $t$-th step and $\alpha_{1},\dots,\alpha_{T}$ is a predefined schedule, where $T$ denotes the total number of steps. An important property brought by the Markov chain is that we can obtain $x_{t}$ from $x_{0}$ directly via:
$\small q({\bm{x}}_{t}|{\bm{x}}_{t-1})=\mathcal{N}\left({\bm{x}}_{t};\sqrt{\alpha_{t}}{\bm{x}}_{t},(1-\alpha_{1})\mathbf{I}\right),$ (6)
@@
DMs use a network $p_{\bm{\theta}}({\bm{x}}_{t-1}|{\bm{x}}_{t},t)$ to fit the real distribution $q({\bm{x}}_{t-1}|{\bm{x}}_{t})$. The overall simplified optimization target is a sampling and denoising process as follows:
$\small L_{\text{simple}}(\bm{\theta})=\mathbb{E}_{t,{\bm{x}}_{0},\bm{\epsilon}}\left[\left|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\sqrt{\alpha_{t}}{\bm{x}}_{0}+\sqrt{1-\alpha_{t}}\bm{\epsilon},t)\right|^{2}\right].$ (8)
@@
Figure 2: Illustration from the data generation over the transformation to train a binary classifier $C$. Adversarial and benign samples are separately transformed. The transformation implies that the input image $\mathbf{x}_{0}$ is first gradually inverted into a noise image $\mathbf{x}_{T}$ using DDIM inversion [19], and then it is denoised step by step until the transformed $\mathbf{x}_{0}^{\prime}$ is obtained, as illustrated in eq. 13.
@@
$\begin{split}{\bm{x}}_{t-1}&=\sqrt{\alpha_{t-1}}\left(\frac{{\bm{x}}_{t}-\sqrt{1-\alpha_{t}}\bm{\epsilon}_{\bm{\theta}}({\bm{x}}_{t},t)}{\sqrt{\alpha_{t}}}\right)+\\ &\sqrt{1-\alpha{t-1}-\sigma^{2}_{t}}\cdot\bm{\epsilon}_{\bm{\theta}}({\bm{x}}_{t},t)+\sigma_{t}\bm{\epsilon}_{t}.\end{split}$ (9)
@@ 3.2 Method Details
Given an input image ${\bm{x}}_{0}$, our objective is to classify whether it is adversarial or natural (benign). To achieve this, we utilize a pre-trained diffusion model, specifically a DDIM, and apply the inversion process, gradually introducing Gaussian noise (refer to eq. 12). After $T$ steps, ${\bm{x}}_{0}$ transforms into ${\bm{x}}_{T}$, which now belongs to an isotropic Gaussian noise distribution. Subsequently, we apply the reverse process (refer to 9) to convert the noisy image, resulting in a recovered version ${\bm{x}}$. The overall transformation is defined as:
@@
$L(y,y^{\prime})=-\sum_{i=1}^{N}\left(y_{i}\log(y^{\prime}_{i})+(1-y_{i})\log(y^{\prime}_{i})\right),$ (14)
@@ 4.2 Evaluation metrics
3:Generate transformed dataset $\mathcal{X}_{\text{tf}}$ using the diffusion model

### 01 Mar 2024 v4

@@ 1 INTRODUCTION
Figure 1: Illustration of the difference between an adversarial and a benign sample when subjected to the transformation process. $p_{a}({\bm{x}})$ represents the distribution of adversarial images, while $p_{b}({\bm{x}})$ represents the distribution of benign images. Using the inversion and reversion process of a pre-trained DDIM [19], ${\bm{x}}_{a}$ and ${\bm{x}}_{b}$ become ${\bm{x}}_{a}^{\prime}$ and ${\bm{x}}_{b}^{\prime}$, respectively. These transformed counterparts now belong to distinct distributions, namely $p_{a}^{\prime}({\bm{x}})$ and $p_{b}^{\prime}({\bm{x}})$, characterized by a significantly reduced overlap. Therefore, this results in a distinct representation of adversarial samples compared to benign samples.
@@ 3 METHOD
In this section, we begin with reviewing DDPMs, and the inverse and reverse process of the DDIM [19]. Then, we present details of our method, based on [59] originated for deepfake detection, and how to capture the adversarial attack fingerprints.
@@ 3.1 Prelimaries
In the following, we use the notations from DDIM [19] because we use this architecture throughout this paper and is also able to use the pre-trained weights from the DDPM architecture [23]: We note that in [23], a diffusion hyperparameter $\beta_{t}$ is first introduced, and then relevant variables $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{t=1}^{T}\alpha_{t}$ are defined. From the DDIM paper we use the notation $\alpha_{t}$ to represent $\bar{\alpha}_{t}$ and also define
$\beta_{t}=1-\frac{\alpha_{t}}{\alpha_{t-1}}.$ (5)
@@
$\small q({\bm{x}}_{t}|{\bm{x}}_{t-1})=\mathcal{N}\left({\bm{x}}_{t};\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}{\bm{x}}_{t-1},(1-\frac{\alpha_{t}}{\alpha_{t-1}})\mathbf{I}\right),$ (6)
in which ${\bm{x}}_{t}$ denotes the noisy image at the $t$-th step and let $\alpha_{1},\dots,\alpha_{T}\in(0,1]^{T}$ be a predetermined decreasing schedule, where $T$ represents the total number of steps. An essential property conferred by the Markov chain is the direct derivation of $x_{t}$ from $x_{0}$, as follows:
$\small q({\bm{x}}_{t}|{\bm{x}}_{0})=\mathcal{N}\left({\bm{x}}_{t};\sqrt{\alpha_{t}}{\bm{x}}_{0},(1-\alpha_{1})\mathbf{I}\right),$ (7)
@@
DMs employ a network $p_{\bm{\theta}}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ to approximate the real distribution $q({\bm{x}}_{t-1}|{\bm{x}}_{t})$. The primary goal of optimization is to achieve a sampling and denoising process as outlined below:
$\small L_{\text{simple}}(\bm{\theta})=\mathbb{E}_{t,{\bm{x}}_{0},\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\sqrt{\alpha_{t}}{\bm{x}}_{0}+\sqrt{1-\alpha_{t}}\bm{\epsilon},t)\right\|^{2}\right].$ (9)
@@
Figure 2: Illustration from the data generation over the transformation through a pre-trained DDIM to train a binary classifier $C$. Adversarial and benign samples are separately transformed. The transformation implies that the input image $\mathbf{x}_{0}$ is first gradually inverted into a noise image $\mathbf{x}_{T}$ using DDIM inversion [19], and then it is denoised step by step until the transformed $\mathbf{x}_{0}^{\prime}$ is obtained, as illustrated in eq. 14.
@@
$\begin{split}{\bm{x}}_{t-1}&=\sqrt{\alpha_{t-1}}\left(\frac{{\bm{x}}_{t}-\sqrt{1-\alpha_{t}}\bm{\epsilon}_{\bm{\theta}}({\bm{x}}_{t},t)}{\sqrt{\alpha_{t}}}\right)+\\ &\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\cdot\bm{\epsilon}_{\bm{\theta}}({\bm{x}}_{t},t)+\sigma_{t}\bm{\epsilon}_{t}.\end{split}$ (10)
@@ 3.2 Method Details
Given an input image ${\bm{x}}_{0}$, our objective is to classify whether it is adversarial or natural (benign). To achieve this, we utilize a pre-trained diffusion model, specifically a DDIM, and apply the inversion process, gradually introducing Gaussian noise (refer to eq. 13). After $T$ steps, ${\bm{x}}_{0}$ transforms into ${\bm{x}}_{T}$, which now belongs to an isotropic Gaussian noise distribution. Subsequently, we apply the reverse process (refer to 10) to convert the noisy image, resulting in a recovered version ${\bm{x}}$. The overall transformation is defined as:
@@
$\ell(y,y^{\prime})=-\sum_{i=1}^{N}\left(y_{i}\log(y^{\prime}_{i})+(1-y_{i})\log(y^{\prime}_{i})\right),$ (15)
@@ 4.2 Evaluation Metrics
3:Generate transformed dataset $\mathcal{X}_{\text{tf}}$ using a pre-trained diffusion model